OpenStudy (anonymous):

One CD player is 3 times as much as another. Together the two CD players would be $320. What is the price of the more expensive CD player?

OpenStudy (anonymous):

Let x be the more expensive CD player and y be the less expensive one. You can make a system of two equations: x = 3y x + y = 320 Solve by substitution: x + y = 320 (substitute 3y for x) 3y + y = 320 (combine like terms) 4y = 320 (divide both sides by 4) y = 80 x = 3y x = 3(80) x = 240 ANSWER: The price of the more expensive CD player is $240.

OpenStudy (anonymous):

hope it helps!! :D

OpenStudy (anonymous):

what grade is this stuff?not being rude just asking